# Denoising Functional Maps: Diffusion Models for Shape Correspondence

### v1

@@ 5.2 Near-Isometric Shape Matching
  • The FAUST dataset contains meshes of 10 humans with various body types in 10 poses, 100 in total. The training/testing split is 80/20.
@@
To compare both categories of baselines fairly, we note that SCAPE is too small to train large models. On the other hand, FAUST was created with SMPL [37] like SURREAL, and includes a sufficient diversity of body types and poses. Therefore, we follow [11, 69, 10] and train large models on SURREAL, descriptor-based models on FAUST, and evaluate them on three test datasets. A comparison with the baselines trained on both FAUST and SCAPE can be found in Supplement C.3.
@@
Figure 3: Qualitative results by our approach on SHREC’19 (top) and DT4D (bottom) datasets; see also Supplement C.2.
@@ Results.
We report the mean geodesic error of our model and the baselines in Table 1. Compared to other large-scale methods trained on SURREAL, our work generalizes well to meshes from SCAPE and SHREC’19. A common problem in large-scale baselines is predicted point maps with incorrect symmetry. This is caused by the fact that the geometry of the evaluation meshes has fewer fine-scale details such as fingers, which help to disambiguate the internal symmetry. Our work avoids overfitting to fine-scale details of SURREAL by using data augmentation (Sec. 5.1). As a result, we can utilize large amounts of data more effectively than previous works.
@@ 5.3 Matching with Anisotropic Meshing
Table 2: Results on the anisotropic versions of FAUST and SCAPE. Large-scale methods are trained on SURREAL, and descriptor-based methods on FAUST.
@@ 5.4 Zero-Shot Shape Matching
Generalization to shape types not observed during training is a desirable property in many shape matching applications. To evaluate our model on unseen data, we use the non-isometric DT4D dataset [40] that consists of nine humanoid shape classes. While typically, 198 shapes are defined for training and 95 for testing, we instead make the evaluation more challenging by training on human datasets, SURREAL or FAUST. This can be viewed as a zero-shot setting, since most of the test shapes are different from the training data (see Supplement C.2 for an illustration).
@@
Table 3: Results on shape classes from DT4D [40]. The best results are highlighted, where applicable. Classes mousey and ortiz are not present in inter-class split. All inter-class evaluation pairs have the format “crypto vs other class”, so crypto is also excluded.
@@ Results.
Our proposed experimental setup is challenging due to the differences between the training and evaluation shapes. Therefore, we use a lower-resolution model predicting a 32$\times$32 functional map, results of all our models are shown in Supplement C.6. We compare our method with recent unsupervised descriptor-based approaches: ConsistentFMaps [69] and ULRSSM [11].
The mean geodesic errors per class for inter- and intra-category splits are shown in Table 3, additional comparison can be found in Supplement C.8. Our method achieves accurate matching for most shape classes in both splits, demonstrating the ability of our method to generalize to data unseen during training. We show the qualitative examples in Fig. 3.
@@ 5.5 Application to Animal Shapes
To show the generality of our method to shape classes other than humans, we test our approach on the SMAL dataset containing seven animal types. We take the 32/17 split from [16], where all categories are present in train/test sets. While the original training set is sufficient for descriptor-based baselines, it is too small for our model. To address this, we use the fitted parameters of the SMAL parametric model [77] and randomly vary the pose of each training animal, generating a total of 64,000 meshes. We use the mean SMAL shape as a template. Note that this generated dataset conveys little new information, since e.g., there are only three different types of lions.
@@ Results.
Table 4: Results on the SMAL dataset [16], all categories are present in train/test sets. We additionally show separate results for each category, e.g. lions vs. everything.
@@ Acknowledgement.
This supplementary material presents several important components of our work that were not included in the main text:
@@ Appendix A Sign Correction
This section provides a quantitative and qualitative evaluation of the sign correction network. This component of our model allows to select a specific sign for eigenvectors of the Laplacian by projecting them onto the learned correction vector and making the projection positive (Sec. 4.2). After performing the sign correction on all eigenvectors, we obtain a specific basis, independent of the initial one provided by the numerical eigensolver.
@@ A.1 Accuracy
(a) Learned correction vector $\varsigma_{24}$ (left) and the corresponding eigenvector $\phi_{24}$ (right). The projection is $0.51$, so we keep the sign of the eigenvector.
@@
Figure 5: Learned correction vectors and their corresponding eigenvectors. Positive and negative values are shown in red and blue, respectively. The low-order correction vectors resemble the eigenvectors themselves, while the high-order ones are mainly concentrated in the arm and hand regions.
(b) Examples on the DT4D intra-category dataset, compared to ULRSSM [11], a state-of-the-art descriptor-based approach. Our method outperforms it on most shape classes, except for a few categories that are too different from the training data: “Mousey”, “Skeletonzombie”, “Ortiz”. The correspondences for the “Drake” class are mostly correct but sometimes have flipped left-right symmetry.
Figure 6: Additional qualitative examples. Our method provides correct correspondences for many challenging human and humanoid meshes.
@@ A.2 Distribution of Functional Maps
Here we study the change in the distribution of functional maps after sign correction. We use the FAUST [7] dataset, which contains meshes of 10 classes of humans with various body types in different poses. For each mesh, we obtain a 32-dimensional eigenbasis and correct it with a pretrained sign correction network. We then compute the template-wise functional maps, as well as the conditioning matrix $y$ (Eq. 8).
Next, we perform PCA decomposition [2] on the functional maps before and after sign correction, as well as on the conditioning matrix. The pairwise plots of the first two components are shown in Fig. 4, where each point is colored according to its class. As we can see, before sign correction, the functional maps did not have a sensible distribution since each of them was defined in a random basis. After the sign ambiguity was resolved, the functional maps form noticeable clusters grouped by class. For the conditioning matrix, the overall structure resembles the distribution of functional maps after sign correction. This explains the basic principle of our method: the sign correction network transformed the space of functional maps into a learnable distribution by making them refer to a specific basis, and the denoising diffusion model approximated this distribution.
(a) FAUST
(b) SCAPE
(c) SHREC’19
Figure 9: Limitations of our method. They include shapes with large differences from the synthetic human data, as well as topological inconsistencies like missing body parts. These limitations may be addressed in future work.
A.3 Learned Correction Vectors
We visualize several of the learned correction vectors in Fig. 5, along with the corresponding eigenvectors. As we can see, the low-order correction vectors resemble the eigenvectors themselves, while the high-order ones are mostly concentrated in the arm and hand regions.
@@ B.1 Robustness to Basis Ambiguity
A straightforward way to account for the basis ambiguity is to use a single correction vector $\varsigma_{i}$ for several adjacent eigenvectors instead of only one eigenvector $\phi_{i}$. In practice, we split the eigenvectors into groups of $32$ and select increasingly more eigenvectors per correction vector for each group: one for $\phi_{1}-\phi_{32}$, two for $\phi_{33}-\phi_{64}$, and two or four for $\phi_{65}-\phi_{96}$. This takes into account the fact that high-order eigenvalues are more likely to have high multiplicity.
@@ Appendix C Additional Evaluation
C.1 PCK Curves
C.2 Qualitative Examples
C.3 Baselines Trained on FAUST and SCAPE
In Sec. 5.2, we evaluated descriptor-based methods that were trained on the FAUST dataset since it contains diverse SMPL-like shapes, while SCAPE is too small to fine-tune large models. For a complete comparison, here we additionally evaluate baselines trained on both FAUST and SCAPE. The results are shown in Table 6. As we can see, our method still performs on par with state-of-the-art methods, even without additional variety provided by the shapes from SCAPE.
@@ C.3 Baselines Trained on FAUST and SCAPE
C.4 Baselines Trained on SURREAL
@@
C.5 Impact of ZoomOut
To illustrate the role that ZoomOut refinement [44] plays in our pipeline, we tested our models without the spectral upsampling step. The results are shown in Tables 8(a) and 8(b). As we can see, ZoomOut plays a key role for 32$\times$32 and 64$\times$64 models, while it is less important for the 96$\times$96 model.
@@ C.5 Impact of ZoomOut
C.6 DT4D: Impact of the Model Dimension
@@ C.7 Ablation Study
C.8 Comparison with ULRSSM on DT4D
In Sec. 5.4, we evaluated our method on shape classes from the DT4D dataset against the baselines, ConsistentFMaps [69] and ULRSSM [11]. Compared to other descriptor-based methods, ULRSSM fine-tunes the parameters of the feature extractor by backpropagating the unsupervised loss during test time for each evaluation pair. For a complete comparison, we evaluate ULRSSM without refinement, which leads to worse results of this baseline; see Table 11.

### v2

@@ 5.2 Near-Isometric Shape Matching
  • The FAUST dataset contains meshes of 10 human classes in 10 poses, 100 in total. The train/test split is 80/20.
@@
To compare both categories of baselines fairly, we note that SCAPE is too small to train large models. On the other hand, FAUST was created with SMPL [37] like SURREAL, and includes a sufficient diversity of body types and poses. Therefore, we follow [11, 69, 10] and train large models on SURREAL, descriptor-based models on FAUST, and evaluate them on three test datasets. A comparison with the baselines trained on FAUST+SCAPE can be found in Supp. C.2.
@@
Figure 3: Qualitative results by our approach on SHREC’19 (top) and DT4D (bottom) datasets; see also Supp. C.1.
@@ Results.
We report the mean geodesic error of our model and the baselines in Table 1. Compared to other large-scale methods trained on SURREAL, our work generalizes well to meshes from SCAPE and SHREC’19. A common problem in large-scale baselines is predicted point maps with incorrect symmetry. This is caused by the fact that the geometry of the evaluation meshes has fewer fine-scale details such as fingers, which help to disambiguate the internal symmetry. Our work avoids overfitting to fine-scale details of SURREAL by using data augmentation (Sec. 5.1). As a result, we can utilize large amounts of data more effectively.
@@ 5.3 Matching with Anisotropic Meshing
Table 2: Results on the anisotropic FAUST and SCAPE.
@@ 5.4 Zero-Shot Shape Matching
Generalization to shape types not observed during training is a desirable property in many shape matching applications. To evaluate our model on unseen data, we use the non-isometric DT4D dataset [40] that consists of nine humanoid shape classes. While typically, 198 shapes are defined for training and 95 for testing, we instead make the evaluation more challenging by training on human datasets, SURREAL or FAUST. This can be viewed as a zero-shot setting, since most of the test shapes are different from the training data (see Supp. C.1 for an illustration).
@@
Table 3: Results on shape classes from DT4D. Classes mousey and ortiz are not present in inter-class split. All inter-class evaluation pairs have the format “crypto vs other class", so crypto is excluded.
@@ Results.
Our proposed experimental setup is challenging due to the differences between the training and test shapes. We use a lower-resolution model predicting a 32$\times$32 functional map, results of all our models are shown in Supp. C.5. We compare our method with recent descriptor-based approaches: ConsistentFMaps [69] and ULRSSM [11].
The mean geodesic errors per class for inter- and intra-category splits are shown in Table 3, with further comparisons in Supp. C.7. Our method achieves accurate matching for most shape classes in both splits, demonstrating the generalization of our method to data unseen during training. We show the qualitative examples in Fig. 3.
@@ 5.5 Application to Animal Shapes
To demonstrate the applicability of our method to shape classes other than humans, we test our approach on the SMAL dataset containing seven animal types. We take the 32/17 split from [16], where all categories are present in train/test sets. While the original training set is sufficient for descriptor-based baselines, it is too small for our model. To address this, we use the fitted parameters of the SMAL parametric model [77] and randomly vary the pose of each training animal, generating a total of 64,000 meshes. We use the mean SMAL shape as a template. Note that this generated dataset conveys little new information, since e.g., there are only three different types of lions.
@@ Results.
Table 4: Results on the SMAL dataset [16], all categories are present in train/test sets. Format: category vs. everything.
@@ Acknowledgement
This supplementary material presents several components of our work that were not included in the main text:
@@ Appendix A Sign Correction
This section provides a quantitative and qualitative evaluation of the sign correction network. This component of our model allows to select a specific sign for eigenvectors of the Laplacian by projecting them onto the learned feature vectors and making the projection positive (Sec. 4.2). After performing the sign correction on all eigenvectors, we obtain a specific basis, independent of the initial one provided by the numerical eigensolver.
@@ A.1 Accuracy
(a) Learned feature vector $\varsigma_{24}$ (left) and the corresponding eigenvector $\phi_{24}$ (right). The projection is $0.51$, so we keep the sign of the eigenvector.
@@
Figure 5: Learned feature vectors and their corresponding eigenvectors. Positive and negative values are shown in red and blue, respectively. The low-order feature vectors resemble the eigenvectors themselves, while the high-order ones are mainly concentrated in the arm and hand regions.
@@ A.2 Distribution of Functional Maps
Here we study the change in the distribution of functional maps after sign correction. We use the FAUST [7] dataset, which contains meshes of 10 classes of humans with various body types in different poses. For each mesh, we obtain a 32-dimensional eigenbasis and correct it with a pretrained sign correction network. We then compute the template-wise functional map and the conditioning matrix $y$ (Eq. 7).
Next, we perform PCA decomposition [2] on the functional maps before and after sign correction, as well as on the conditioning matrix. The pairwise plots of the first two components are shown in Fig. 4, where each point is colored according to its class. As we can see, before sign correction, the functional maps did not have a sensible distribution since each of them was defined in a random basis. After resolving the sign ambiguity, the functional maps form noticeable clusters grouped by class. For the conditioning matrix, the overall structure resembles the distribution of functional maps after sign correction. This explains the basic principle of our method: the sign correction network transformed the space of functional maps into a learnable distribution by making them refer to a specific basis, and the denoising diffusion model approximated this distribution.
A.3 Learned Features
We visualize several of the learned feature vectors in Fig. 5, along with the corresponding eigenvectors. As we can see, the low-order feature vectors resemble the eigenvectors themselves, while the high-order ones are mostly concentrated in the arm and hand regions.
@@ B.1 Robustness to Basis Ambiguity
A straightforward way to account for the basis ambiguity is to use a single feature vector $\varsigma_{i}$ for several adjacent eigenvectors $\left(\phi_{i},\phi_{i+1}\ldots\right)$ instead of only one eigenvector $\phi_{i}$. In practice, we split the eigenvectors into groups of $32$ and select increasingly more eigenvectors per feature vector for each group: one for $\phi_{1}-\phi_{32}$, two for $\phi_{33}-\phi_{64}$, and four for $\phi_{65}-\phi_{96}$. This takes into account the fact that high-order eigenvalues are more likely to have high multiplicity.
@@ Appendix C Additional Evaluation
C.1 Qualitative Examples
(b) Examples on the DT4D intra-category dataset, compared to ULRSSM [11] and ConsistentFMaps [69], state-of-the-art descriptor-based approaches. Our method outperforms them on most shape classes, except for a few categories that are too different from the training data: “Mousey”, “Skeletonzombie”, “Ortiz”. The correspondences for the “Drake" class are mostly correct but sometimes have flipped left-right symmetry.
Figure 7: Additional qualitative examples. Our method provides correct correspondences for challenging human and humanoid meshes.
C.2 Baselines Trained on FAUST and SCAPE
Previously we evaluated descriptor-based methods that were trained on the FAUST dataset since it contains diverse SMPL-like shapes, while SCAPE is too small to fine-tune large models (Sec. 5.2). For a complete comparison, here we additionally evaluate baselines trained on both FAUST and SCAPE. The results are shown in Table 6. As we can see, our method still performs on par with state-of-the-art methods, even without additional variety provided by the shapes from SCAPE.
@@ C.2 Baselines Trained on FAUST and SCAPE
C.3 Baselines Trained on SURREAL
@@
C.4 Impact of ZoomOut
To illustrate the role that ZoomOut refinement [44] plays in our pipeline, we tested our models without the spectral upsampling step. The results are shown in Tables 8(a)-8(b). As we can see, ZoomOut plays a key role for 32$\times$32 and 64$\times$64 models, while it is less important for the 96$\times$96 one.
C.5 DT4D: Impact of the Model Dimension
@@ C.6 Ablation Study
(a) FAUST
(b) SCAPE
(c) SHREC’19
Figure 9: Limitations of our method. They include shapes with large differences from the synthetic human data, as well as topological inconsistencies like missing body parts. These limitations can be addressed in future work.
@@
C.7 Comparison with ULRSSM on DT4D
In Sec. 5.4, we evaluated our method on shape classes from the DT4D dataset against the baselines, ConsistentFMaps [69] and ULRSSM [11]. Compared to other descriptor-based methods, ULRSSM fine-tunes the parameters of the feature extractor by backpropagating the unsupervised loss during test time for each evaluation pair. For a complete comparison, in Table 11 we evaluate ULRSSM without refinement, which leads to worse results of this baseline.